You placed $7,527 in a savings account today that earns an annual interest rate of 10.25 percent, compounded semiannually. How much will you have in this account at the end of 2 years? Assume that all interest received at the end of the period is reinvested the next period. Round the answer to two decimal places.
Answer #1

P=Amount Invested today=$7527

Annual Interest Rate=10.25%=0.1025

i= Semiannual interest rate=0.1025/2=0.05125

N=Number of semi annual periods=2*2=4

CAF=Compound Amount Factor(F/P,i,N)=(1+i)^N

i=0.05125, N=4

CAF=(F/P,0.05125,4)=(1+0.05125)^4=

1.221305

F=Amount in Account after 2 years=7527*1.221305=

$ 9,192.76

Amount in account after 2 years $ 9,192.76